### v1

# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

Abstract
The realization of universal robots is an ultimate goal of researchers. However, a key hurdle in achieving this goal lies in the robots’ ability to manipulate objects in their unstructured surrounding environments according to different tasks. The learning-based approach is considered an effective way to address generalization. The impressive performance of foundation models in the fields of computer vision and natural language suggests the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we believe achieving general manipulation capability requires an overarching framework akin to auto driving. This framework should encompass multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability. This survey focuses on the contributions of foundation models to robot learning for manipulation. We propose a comprehensive framework and detail how foundation models can address challenges in each module of the framework. What’s more, we examine current approaches, outline challenges, suggest future research directions, and identify potential risks associated with integrating foundation models into this domain.
keywords:
Robot learning, manipulation, foundation model, survey, universal robot
1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ”I, Robot”. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured surrounding environments according to different tasks. There is abundant literature available for improving the general manipulation capabilities of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
The predominant methodologies in learning-based approaches are deep learning, reinforcement learning and imitation learning. Learning-based methods have spanned from acquiring specific manipulation skills through labeled datasets like human demonstration, to acquiring abstract representations of manipulation tasks conducive to high-level planning, to exploring an object’s functionalities through interaction and encompassing various objectives in between (Kroemer et al. (2021)). However, challenges persist, including 1) unnatural interaction with humans; 2) high-cost data collection; 3) limited perceptual capabilities; 4) non-intelligent hierarchy of skills; 5) inaccurate pre- and post-conditions & post-hoc correction; 6) unreliable skill learning; 7) poor environment transition (Hu et al. (2023b)).
Foundation models are primarily pretrained on vast internet-scale datasets, enabling them to be fine-tuned for diverse tasks. Their significant advancements in vision and language processing contribute to mitigating the aforementioned challenges. Based on the different input modalities and functionalities of the models, we categorize foundation models into the following six types.
  1. 1.
    Large Language Models (LLMs) like BERT (Devlin et al. (2018)), GPT-3 (Brown et al. (2020)) have demonstrated the capability to generate coherent chains of thought.
  2. 2.
    Vision Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capabilities for open-set objects.
  3. 3.
    Visual Content Generation Models (VGMs) like DALL-E (Ramesh et al. (2021)), Zero-1-to-3 (Liu et al. (2023c)), demonstrate the capability to generate 2D images or 3D meshes through text or images.
  4. 4.
    Vision Language Models (VLMs) like GPT-4V (Achiam et al. (2023)), CLIP (Radford et al. (2021)) showcase robust comprehension of both vision and language, such as open-set image classification and visual question answering.
  5. 5.
    Large Multimodal Models (LMMs) expand their scope beyond vision and language to create novel categories of foundation models incorporating additional modalities, such as ULIP (Xue et al. (2023a)) aligns point cloud representation to the pre-aligned image-text feature space.
  6. 6.
    Robot Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observation-to-action mapping.
In this survey, we investigate how foundation models are utilized in robot learning for manipulation, like Fig. 1:
  1. 1.
    LLMs enable the direct generation of policy codes or action sequences and facilitate natural interaction with the environment.
  2. 2.
    VFMs enhance open-world perception.
  3. 3.
    VLMs serve as the cornerstone for alignment between vision and language, facilitating understanding of multimodality.
  4. 4.
    LMMs expand their modalities to include 3D point cloud and haptic data, among others.
  5. 5.
    VGMs generate 2D images or 3D meshes based on prompting, aiding in scene generation within simulation environments.
  6. 6.
    RFMs serve as an end-to-end policy model, directly outputting actions based on input observations
Figure 1: LLMs help address challenges in Interaction, Manipulation Data Generation, Hierarchy of Skills, Skill Policy Learning, and Environment Transition Model. VLMs assist in tackling challenges in Interaction, Manipulation Data Generation, Hierarchy of Skills, Pre- and Post-conditions Detection, Skill Policy Learning, and Perception. LMMs aid in addressing challenges in Interaction and Perception. VGMs tackle the challenge of Manipulation Data Generation. VFMs help address challenges in Manipulation Data Generation, Hierarchy of Skills, Pre- and Post-conditions Detection, Skill Policy Learning, and Perception. RFMs assist in addressing the challenge of Skill Policy Learning.
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capabilities. However, we do not believe that a single foundation model alone can achieve general manipulation capabilities. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% safety rate in manipulation tasks, remains a challenge. Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capabilities necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capabilities.
The ultimate general manipulation framework should be capable of interacting with humans to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation objectives. However, the interaction between agent and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios can be static or dynamic, with objects being either rigid or deformable and task objectives ranging from short-horizon to long-horizon. We designate the restriction of the agent’s learning capability to improving old skills and manipulating rigid objects in static scenes to achieve short-horizon task objectives as Level 0 (L0) . Drawing from Kroemer et al. (2021), we propose a comprehensive framework for general manipulation task at L0. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
Di Palo et al. (2023) and Firoozi et al. (2023) provide detailed descriptions of the application of foundation models in navigation and manipulation, but these lack thoughtful consideration of the relationship between foundation models across different applications. The survey most closely related to this paper is Xiao et al. (2023). Compared to this survey, our survey focuses on the contributions of foundation models to robot learning for manipulation, proposing a comprehensive framework and detailing how foundation models can address challenges in each module of the framework.
This paper is structured as follows: In Sec. 2, we present a comprehensive framework of robot learning for general manipulation, based on the developmental history of robot learning. We elaborate on the impact of foundation models on each module in the framework in the following sections. Sec. 3 is Human/Agnet Interaction module, Sec. 4 is Pre- and Post-conditions Detection module, Sec. 5 is Hierarchy of Skills module, Sec. 6 is State Perception module, Sec. 7 is Policy module, Sec. 8 is Manipulation Data Generation module. In Sec. 9, we discuss several issues of particular concern to us. In Sec. 10, we summarize the contributions of this survey and identify the limitations of the current framework as well as the ongoing challenges.
2 Framework of Robot Learning for General Manipulation
Over the past decade, there has been a significant expansion in research concerning robot manipulation, with a focus on leveraging the growing accessibility of cost-effective robot arms and grippers to enable robots to interact directly with the environment in pursuit of their objectives. As the real world encompasses extensive variation, a robot cannot expect to possess an accurate model of its surrounding environment, the objects within it, or the skills necessary for manipulation in advance (Kroemer et al. (2021)).
Early stage, robotic manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023)) and Graspnet (Fang et al. (2020)) take point cloud as input and output the target poses.
In the mid-term, many tasks in robotics require a series of correct actions, which are often long-horizon tasks. For example, making a cup of tea with a robot involves multiple sequential steps such as boiling water, adding a tea bag, pouring hot water, etc. Learning to plan for long horizons is a central challenge in episodic learning problems (Wang et al. (2020)). Decomposing tasks has several advantages. It makes learning individual skills more efficient by breaking them into shorter-horizon, thus aiding exploration. Reusing skills in multiple settings can speed up learning by avoiding the need to relearn elements from scratch each time. Researchers train a hierarchy model $H$ to decompose the task into a sequence of subgoals (Ahn et al. (2022)), and observe pre- and post-conditions to ensure that the prerequisites and outcomes of each subgoals are met (Cui et al. (2022)). This comes from human experience in task completion, breaking tasks down into several steps and being able to decide on the execution method for each step based on the initial environment, such as putting a tea bag into a teacup using either a pick-place or pushing method (pre-condition), and ultimately detecting whether the step is completed to proceed to the next step (post-condition).
Recently, researchers have realized that training policies require real-world interaction between robots and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, agents can generate samples based on it, significantly reducing the frequency of direct interaction between agents and environments (Liu et al. (2024)).
We organize the framework of robot learning for general manipulation according to its development history, as shown in Fig. 2. Due to the task instruction and corrective instruction are commanded by humans or other agents, we include an interaction module $I$ in the framework. The aforementioned modules all require datasets for learning. The data collection process for the hierarchy of skills $H$ and pre- and post-conditions detection modules $P$ is similar to that in the fields of CV and NLP. Compared to data collection in CV and NLP domains, gathering datasets for manipulation tasks involves interaction between robots and the real environment, leading to high costs and risks. A simulation environment can mitigate the challenges of dataset collection for manipulation tasks, but there still exists a sim-to-real gap. Therefore, we emphasize the manipulation data generation $D$, to provide dataset for the state $S$, policy $P$, and transition $T$ modules.
  1. 1.
    Human/Agent Interaction. As a general-purpose robot, it is necessary to understand the task goal instruction conveyed by humans or collaborative agents to accomplish tasks. It is also important to synchronize and receive corrective instructions when one’s actions deviate from the goal or when the goal changes during task execution.
  2. 2.
    Pre- and Post-conditions Detection. Before the hierarchy of skills and executing a skill, check the environment to ensure it meets the requirements and select the execution method based on object affordances. For instance, when placing a tea bag in a teacup, pre-conditions detection determines whether there are tea bags and teacups and chooses between pick-place or pushing based on their affordances, such as, a tea bag is spherical, and it has the affordance of rolling when pushed. After executing a skill, verify if the objective is achieved. If not, post-conditions detection identifies the reason for failure. For example, if a teacup is knocked over during pick-and-place, inform post-hoc correction and use pick-and-place to upright the cup and reinsert the tea bag.
  3. 3.
    Hierarchy of skills. The concept of ’Hierarchy of skills’ often involves creating a sequence of subgoals (Song et al. (2023)). Each subgoal necessitates a skill, which may consist of one or multiple primitive actions (Zhang et al. (2023c)). For instance, tasks like filling the kettle with water, heating the water, and getting the tea leaves are examples of subgoals that agents need to achieve in a specific order to fulfill the final goal as instructed.
  4. 4.
    State. The state consists of robot proprioception $S_{robot}$, general environmental state $S_{e}$, and object states $S_{o}$. The difference between $S_{e}$ and $S_{o}$ is analogous to the foreground and background of an image. $S_{robot}$ generally relates to the mechanical structure of robots. Currently, there are limited studies focusing on the improvement of robot mechanical structures using foundation models, with Stella et al. (2023) being one of them. However, researches in this direction are scarce and still in their initial stages.
  5. 5.
    Policy. The policy outputs actions to accomplish task goals based on the input state. At presents, the paradigms for generating actions using foundation models include LLMs directly generating code for robot execution, VLMs directly generating or VLMs combined with LLMs generating corresponding target poses, RFMs directly outputting target pose or delta pose through end-to-end training, and foundation models assisting reinforcement learning in generating corresponding actions.
  6. 6.
    Action. Different foundation models applied as policies result in different types of actions. We categorize them into three types: Code, Target Pose, and Delta Pose. Code refers to the direct control code of the robot arm and gripper. Target Pose refers to the desired pose of the end-effector, which is input to motion planning to generate the trajectory. Delta Pose refers to the next waypoint the end-effector moves to, with continuously outputted delta poses forming the trajectory.
  7. 7.
    Transition. The transition model can directly output the changed state of the environment and objects based on the input action, thus helping to reduce the interaction between the robot and the real environment. Zhang and Soh (2023) uses LLM as a human model, predicting the changes in human state after robot actions, and selecting the most suitable robot actions for interaction with humans. In various algorithms, the transition environment model is not a mandatory module. Therefore, in Fig. 2, we enclose this module with a dashed box.
  8. 8.
    Manipulation Data Generation. The manipulation data generation module provides manipulation of task-level datasets.
Figure 2: Framework of Robot Learning for General Manipulation. Human or other agents convey task instruction or corrective instruction to the Hierarchy of Skills module $H$ through the Interaction module $I$. Pre-conditions Detection module $P$ conveys object detection and affordance detection perception information to Hierarchy of Skills $H$. Hierarchy of Skills may encounter ambiguity in generating subgoals, such as when there are two cups in the environment, it needs to determine which cup to pour water into through interaction with the external world. After generating subgoals, Policy module $\Pi$ trains according to the dataset generated by Manipulation Data Generation module $D$. During Skill Execution, Action $\alpha$ is generated by perceiving State $S$. Environment Transition $T$ is not a necessary module; its addition can be directly transitioned to the next state through the generated action. When issues arise during execution, Corrective Instruction is conveyed to Policy module $\Pi$ from the external world. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next step; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.
3 Human/Agent Interaction
There are two ways for human or other agent to interact with robot: 1) Providing task instruction to the robot to help it understand the task objective and complete the task independently (Khan et al. (2023)). 2) Collaborating with the human or other agent to complete tasks, sharing workspace information, and conveying corrective instructions when useful or error-correcting information is identified to optimize the robot’s current action (Lynch et al. (2023)).
Figure 3: Interaction Module. Interaction mainly involves the exchange between task instruction and corrective instruction. Ambiguity often arises in task instruction interaction, hence agents need to detect ambiguities. 1) One approach is to perceive objects in a multi-modal environment and enumerate possible interpretations based on perception results (Mo et al. (2023)). 2) Another approach involves the next step prediction module, which predicts and scores the next step; if the scores of the top 2 steps are less than $\delta$, it is considered that the task goal is ambiguous (Ren et al. (2023a)). 3) Strong comprehension skills are required during the transmission of corrective instruction, and the current mainstream approach involves using the encoder of large language models to extract tokens and input them into the policy to modify the original trajectory (Bucker et al. (2023)).
When conveying task instruction to the robot, there may contain language ambiguity in the task goal, such as having both red and green cups in the scene, and the task instruction is ’grasp the cup.’ This ambiguity may confuse the robot regarding which color cup to grasp. To address this issue, the robot needs to inquire about and confirm the final task objective from the human or other agent, thus requiring enhancement of their capabilities in text generation and comprehension. When conveying corrective instruction to a robot, it needs to comprehend the meaning of the corrective instruction and translate corrective instruction into appropriate actions. If needed, the robot should be able to convey its current state to human or agent. For instance, if a robot is picking up a book from a shelf filled with books, lifting too quickly may cause other books to fall. Human or collaborating agent need to alert the robot that the current lifting action is dangerous and advise it to lift slowly. If necessary, the robot should also report its current execution state, such as its grasping speed, and inquire whether this speed is considered high. However, corrective instructions are diverse; thus, understanding them is essential.
In addressing instruction ambiguity and text generation and comprehension challenges, SeeAsk (Mo et al. (2023)) utilizes CLIP’s perceptual module to identify objects in the scene and employs a fixed questioning template to organize language to ask about which object will be manipulated. Although the use of CLIP enhances the generalization ability for object recognition, It can’t generate text for asking questions and to comprehend answers from the outside world and SeeAsk (Mo et al. (2023)) focuses solely on addressing ambiguities concerning object color and spatial relationship due to a fixed questioning template. KNOWNO (Ren et al. (2023a)) utilizes LLM to score the next action to be taken. If the score difference between the top two actions is less than a threshold, it’s considered ambiguous, prompting a confirmation for the final action. This approach improves efficiency and autonomy. Matcha (Zhao et al. (2023b)) not only employs vision but also utilizes haptic and sound senses to perceive object properties, such as material. When encountering ambiguity in object attribute recognition, it leverages LLM to generate inquiry content. CoELA (Zhang et al. (2023b)) utilizes LLM as both a communication module and a planning module to enhance interaction text generation and comprehension, as well as task scheduling, with collaborative agent. LLM-GROP (Ding et al. (2023)) utilizes LLM to extract latent commonsense knowledge embedded within task instructions. For example, a task instruction might be ”set dinner table with plate and fork,” while the latent commonsense knowledge could be ”fork is on the left of a bread plate.”
As for corrective instruction, LILAC (Cui et al. (2023)) utilizes GPT-3 to distinguish between task instructions and corrective instructions. It then employs Distil-RoBERTa to extract text features and input them into the network to modify the robot’s original trajectory. LATTE (Bucker et al. (2023)), on the other hand, employs BERT and CLIP to extract features from corrective instructions and observation images and input them into the network to modify the robot’s original trajectory.
Following Fig. 3, the ability of LLMs to harness thought chains efficiently identifies ambiguity, surpassing the limitations of enumerating ambiguity. LLMs’ comprehension of text effectively understands corrective instructions and transforms the original trajectory into a corrective trajectory.
4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. Object recognition will be addressed in the section 6. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the post-task positions of manipulated objects, yet task failure determination relies on the robotic arm’s execution, such as collision between robot and objects occurrence. Therefore, this section focuses on articles discussing foundation models in object affordance.
4.1 Affordance
The affordances associated with an object represent the range of manipulations that the object affords an agent (Gibson (2014)). Early approaches addressed the issue by treating it as a supervised task (Kokic et al. (2017)). However, the process of annotating datasets is laborious and time-consuming, making it impractical to exhaustively cover all geometric information present in real-world environments. Consequently, researchers are exploring the application of reinforcement learning, enabling robots to collect data and train affordance perception modules through continuous exploration (Wu et al. (2021)). Nevertheless, current reinforcement learning methods are trained in simulated environments, leading to a significant sim-to-real gap. To address these challenges, researchers propose training the affordance perception module using videos of human interactions in real-world scenarios (Ye et al. (2023); Bahl et al. (2023)).
For supervised learning methods, GraspGPT (Tang et al. (2023)) utilizes LLM outputs for object class descriptions and task descriptions. Object class descriptions detail the geometric shapes of each part of an object, while task descriptions outline the desired affordances for task execution, such as the types of manipulation actions to be taken. Integrating both components into the task-oriented grasp evaluator enhances the quality of the generated grasp pose. 3DAP (Nguyen et al. (2023)) utilizes the text encoder of LLM for feature extraction. The extracted features from desired affordances text are inputted into both the affordance detection module and pose generation module. This enhances the quality of the predicted affordance map and the generated pose.
In reinforcement learning, ATLA (Ren et al. (2023b)) utilizes GPT-3 to generate language descriptions of tools. These descriptions are then inputted into a pre-trained BERT model to obtain representations. The extracted features are finally fed into the SAC network module. Meta-learning techniques are employed to enhance the learning efficiency for the use of new tools. Xu et al. (2023a) employ CLIP’s text and image encoders to extract features from language instructions and scene images, improving the quality of grasp pose generation in the SAC module.
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023b)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. LAN-grasp (Mirjalili et al. (2023)) inputs human instructions into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
Following Fig. 4, current research focuses on LLM’s zero/few-shot capability in providing part-level prior knowledge of objects through textual descriptions, which aids in generating affordance maps or grasp poses. While utilizing foundation models directly for generating affordance maps or grasp poses is possible, it still exhibits significant instability, and the performance boundaries remain unclear.
Figure 4: Application of Foundation Models in Affordance Grasp. The main approaches of task-oriented grasp are supervised learning and reinforcement learning. Both methods utilize LLM to generate object part-level description and desired affordance description in task instruction, then fuse tokens into the original network through language encoder to output task-oriented grasp pose (Tang et al. (2023); Ren et al. (2023b)). Direct using foundation method utilizes LLM to generate object part-level description and desired affordance description according to task instruction. VLM marks out the part of the object to grasp in the image based on the description (Liu et al. (2023b)).
5 Hierarchy of skills
In the last two decades, research on skill hierarchy has led to powerful domain-independent planners and various real-world applications (Silver et al. (2022)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos, such as VLaMP (Patel et al. (2023)), which trains models to understand human video operations. Currently, the text-based approach dominates the skill hierarchy domain. PDDL, a Lisp-like language (Silver et al. (2022)), is commonly used to address skill hierarchy issues. However, as LLMs excel in natural language tasks and PDDL is not a natural language, researchers are exploring how LLMs can be employed by embodied agents for skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Additionally, various benchmarks such as PlanBench (Valmeekam et al. (2023)) and CALVIN (Mees et al. (2022)) can assess the planning and reasoning capabilities of LLMs.
LLMs possess a notable limitation: they lack practical experience, hindering their utility for decision-making within a specific context, so the output of LLMs often cannot be translated into executable actions for robots. Huang et al. (2022) first use pre-trained causal LLM to break down high-level tasks into logical mid-level action plans. Then, a pre-trained masked LLM is employed to convert mid-level action plans into admissible actions. The pre-trained causal LLM and pre-trained masked LLM can be the same LLM model, which plays different roles depending on various prompts. However, prompts usually require the context of the robot’s capabilities, its current state, and the environment. LLMs are considered ”forgetful” and don’t treat information in the system prompt as absolute. Despite efforts to reinforce task constraints in the objective prompt and extract numerical task contexts from the system prompt, storing them in data structures, errors caused by LLM forgetfulness remain unresolved (Chen and Huang (2023)).
SayCan (Ahn et al. (2022)) scores pre-trained tasks based on prompting and observation images, generating the task sequence with the highest score. Saycan provides a paradigm for generating action sequences using LLM, but there are still some drawbacks: 1) The generated action sequences do not incorporate user preferences. 2) Safety regulations are not adequately addressed. 3) The limitation of the skill library. 4)LLM focuses solely on reasoning when generating action sequences, neglecting feedback on action execution. 5) The limitation of scene grounding. GD (Huang et al. (2023d)) proposes a paradigm to address the aforementioned issues by not only scoring the generated action sequence using LLM but also introducing a grounded function model for scoring the generated action sequence. The grounded function model encompasses token-conditioned robotic functions, such as affordance functions that capture the abilities of a robot based on its embodiment, safety functions, and more. This approach tackles drawbacks by designing grounded functions, avoiding fine-tuning in LLM.
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023c) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023b)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
As for the limitation of scene grounding, LLMs need to inquire about the scene representation to determine the availability, relationship and location of objects. NLMap (Chen et al. (2023a)) proposes an open-vocabulary, queryable semantic representation map built on ViLD and CLIP. This map outputs the pose of related objects based on task instructions, which are then handed over to the LLM for planning. Text2Motion (Lin et al. (2023)) incorporates a geometric value function on top of the value function, enabling the robot to select actions that adhere to geometric constraints based on scene descriptions. Xu et al. (2023b) explore the possibility of teaching robots to creatively utilize tools within scenarios, which involve implicit physical limitations and require long-term planning. VILA (Hu et al. (2023a)) seamlessly incorporates perceptual data into ChatGPT-4V for its reasoning and planning processes, facilitating a deep comprehension of common sense knowledge within the visual domain, encompassing spatial arrangements and object characteristics. PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to enhance its understanding of physical object attributes, such as material. This integration of a physically informed VLM into an interactive framework with a LLM enhances task planning performance in tasks incorporating instructions related to physical object attributes.
The hierarchy of skills possessed by LLMs or VLMs can be applied not only to single agents but also to multiple agents. SMART-LLM (Kannan et al. (2023)) utilizes LLM for the hierarchy of skills and allocates each task to every agent through the task assignment module.
When utilizing LLMs for skill hierarchy, it provides generalization. Regardless of whether the prompting input to LLMs is in natural language or PDDL format (Silver et al. (2022)), the hierarchy of skills possessed by LLMs or VLMs still exhibits instability. Hence, researchers are exploring approaches that integrate LLMs with classical PDDL-based planning methods for the hierarchy of skills. LLM+P (Liu et al. (2023a)) utilize LLMs to translate natural language into PDDL and input into a classical planner for the hierarchy of skills. Xie et al. (2023b) indicate that LLMs exhibit greater efficacy in translation tasks as opposed to planning.
Following Fig. 5, we focus on the application of video in hierarchy of skills. Video data, being internet-based and possessing large-scale attributes, enables robust generalization when learning high-level task operation descriptions mapped from videos. This approach not only aids in hierarchy of skills but also promotes one-shot/few-shot learning, akin to humans focus on specific high-level manipulation semantic information when learning new task. Utilizing LLMs for hierarchy of skills, we emphasize the GD (Huang et al. (2023d)) framework, which circumvents the high cost of LLM fine-tuning. LLMs exhibit powerful translation abilities, thus offering significant utility in assisting classical planner.
Figure 5: Hierarchy of Skills. 1) Utilize human video operation to learn the skill sequence for task execution, decompose the video of the user’s progress so far into observations and human actions through segmentation, and input them along with task instructions into a pre-trained language model to predict the next step (Patel et al. (2023)). 2) LLM scores the skills in the skill library based on task instructions and the skills already executed, and the value function also scores the skills in the skill library based on observation images. The highest-scoring skill, obtained by multiplying the two scores, is selected as the next step (Ahn et al. (2022)). The value function can consider multiple factors such as affordance, safety, user preference, and more (Huang et al. (2023d)), and these considerations can also be fine-tuning LLM (Wu et al. (2023b)). 3) LLM assists the classical planner by translating task instructions into PDDL descriptions, sending them to the classical planner to generate a PDDL plan, and then translating the PDDL plan into a natural language plan using LLM (Liu et al. (2023a)).
6 State Perception
Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning. The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capabilities for vision models (Dehghani et al. (2023)).
As for pre-trained visual representations, the algorithms mentioned have various training objectives: for instance, contrastive methods like Vi-PRoM (Caron et al. (2021)), R3M (Nair et al. (2022)), VIP (Ma et al. (2022)), CLIP (Radford et al. (2021)), LIV (Ma et al. (2023a)); distillation-based methods such as DINO (Caron et al. (2021)); or masked autoencoder methods like MAP (Radosavovic et al. (2023)), MAE (He et al. (2022)). The primary datasets utilized comprise the CLIP dataset (Radford et al. (2021)), consisting of 400 million (image, text) pairs sourced from the internet, along with ImageNet (Deng et al. (2009)), Ego4D (Grauman et al. (2022)), and EgoNet (Jing et al. (2023)).
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023b)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough. At the same time, it is important to consider the potential of the diffusion model (Ze et al. (2023)).
The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. In the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
As for segmentation, SAM (Kirillov et al. (2023)) develops a transformer-based architecture and creates the largest segmentation dataset, with over 1 billion masks from 11 million images. The model is adaptable and enables zero-shot transfer to new tasks and image distributions. Fast-SAM (Zhao et al. (2023a)) and Faster-SAM (Zhang et al. (2023a)) aim to improve the training and inference speed of the network by enhancing its network structure. TAM (Yang et al. (2023a)) merges SAM (Yang et al. (2023a)) and XMem (Cheng and Schwing (2022)) for high-performance interactive tracking and segmentation in videos.
As for detection, traditional detection models are usually confined to a narrow range of semantic categories because of the cost and time involved in gathering localized training data within extensive or open-label domains. However, advancements in language encoders and contrastive image-text training enable open-set detection. Researchers integrate language into a closed-set detector to generalize open-set concepts, detecting various classes through language generalization despite being trained solely on existing bounding box annotations, such as OWL-ViT (Minderer et al. (2022)), Grounding-DINO (Liu et al. (2023d)), OVD (Zareian et al. (2021)), ViLD (Gu et al. (2021)), DetCLIP (Yao et al. (2022a)).
Deploying such models in open-set detection presents a significant challenge, primarily because even slight alterations in prompting can greatly impact performance. Fine-tuning can enhance a foundation model’s understanding of prompting. However, foundation models are often over-parameterized, leading to slow training processes. COOP (Zhou et al. (2022)) maps prompting to a set of learnable vectors, which can be optimized through network training. In CLIP-Adapter (Gao et al. (2024)), two extra linear layers are appended after the final layer of either the vision or language backbone to enable efficient few-shot transfer learning through fine-tuning.
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with language features to identify the point cloud category.
As for scene reconstruction, due to the foundation model’s robust open-set detection capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are then integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. 3D-LLM (Hong et al. (2023)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023)), endowing 3D features with semantic information. CLIP-NeRF (Wang et al. (2022)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object textures during rendering.
Following Figure 6, representation learning facilitates the learning of the multimodal encoder, aligning features from various modalities to the text domain, aiding the model’s cognition, akin to how humans learn and think through words. The learning of the encoder promotes open-set perception tasks, such as open-set detection, open-set segmentation, and open-set scene reconstruction. When applied to specific scenarios, models often require fine-tuning. CLIP-Adapter (Gao et al. (2024)) offers a fine-tuning method. However, fine-tuning involves different techniques for different model architectures.
Figure 6: State Perception. The representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation, and open-set scene reconstruction. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)), and LERF (Kerr et al. (2023)) utilizes the CLIP (Radford et al. (2021)).
7 Policy
In the realm of manipulation, policies can generate signals for robots’ execution based on input. The types of outputs from policies encompass code, delta pose, and target pose. Policies generate code for robot execution directly, aiding detailed observation by humans for debugging. Generating delta poses directly through policies allows for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose, which often requires subsequent motion planning.
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), and Vision-Language-Action-Delta-Pose (VLADP) based on different output policies.
7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, for a agent capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capabilities, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. In PROGPROMPT (Singh et al. (2023)), assertions are added to the generated code. When executing assertions, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task.
7.2 VLATP
The utilization of foundation models to generate target pose can be categorized into three approaches: 1) Directly using existing foundation models to output target pose. 2) Training RFMs to output target pose through reinforcement learning. 3) Training RFMs to generate target pose through imitation learning.
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image.
Training via the collection of a large manipulation dataset primarily involves utilizing off-line reinforcement learning. PI-QT-Opt (Lee et al. (2023)) leverages a large-scale, multi-task dataset and employs a model-free off-policy reinforcement learning approach for training. RL@Scale (Herzog et al. (2023a)) provides extensive empirical validation through training on real-world data collected over 24 months from experimentation across a fleet of 23 robots in three office buildings in the waste sorting application. Q-Transformer (Chebotar et al. (2023a)) facilitates training high-capacity sequential architectures on mixed-quality data by applying transformer models to RL.
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) and BC-Z (Jang et al. (2022)) demonstrate the capability of imitation learning in language-conditioned general manipulation. Due to the development of language foundation models, LEO (Huang et al. (2023a)) expands upon the language foundation model by incorporating modalities such as images and 3D point clouds and fine-tune manipulation datasets using the LoRA method, showcasing the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023a)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs.
Figure 7: Foundation Models for Policy. 1) Vision-Language-Action-Code (VLAC) integrates task instruction, pre-written APIs, and example inputs into LLM, generating corresponding execution code. Then, when executing code, environment state feedback is obtained to check if the environment satisfies the pre- and post-conditions of the task (Singh et al. (2023)). 2) Vision-Language-Action-Target-Pose (VLATP) without foundation models fine-turning inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation, and both the manipulated object and object segmentation images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D target pose (Huang et al. (2023b)). 3) Vision-Language-Action-Target-Pose (VLATP) with foundation models fine-tuning inputs task instruction and multimodal perception into a pre-trained model such as LLM or VLM after fine-tuning to output target pose (Huang et al. (2023a)). 4) Vision-Language-Action-Delta-Pose (VLADP) without fine-turning. LLM outputs the description of the object to be manipulated based on task instruction. VLM then outputs the BBox of the corresponding object in RGB. Next, segmentation is performed using VFM. 3D Value Map is updated based on depth information, generating delta pose (Huang et al. (2023c)). 5) Vision-Language-Action-Delta-Pose (VLADP) outputs delta pose and predicted observation image after action execution by inputting task instruction and multimodal perception into a pre-trained model after fine-tuning. The difference between VLADP and VLATP with fine-tuning lies in generating delta poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose. Target pose often requires subsequent motion planning. Outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining the image after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM judges the success of execution to provide criteria for RL rewards (Di Palo et al. (2023)).
7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)). However, RobotNet can only achieve one-model corresponding to one-task.
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Though there’s no evidence indicating the superiority of either input method, intuitively, solely observing through images may compromise stability by omitting proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning and proposes a simple but effective self-improvement process and also demonstrates that predicting both the next action and the hindsight image upon executing that action can enhance the performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
Utilizing pre-trained VLMs (Zhang et al. (2024)) for fine-tuning to construct RFMs is considered efficient. GR-1 (Wu et al. (2023a)) is initially pre-trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates sample data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). Therefore, RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses. Human intervention is added, enabling robots to adjust their trajectories promptly based on human input. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023b)) emerges, effectively enabling a robot manipulation policy with VLMs.
Utilizing pre-trained foundation models without fine-tuning reduces the data collection and training steps. Voxposer (Huang et al. (2023c)) utilizes RGB-D observations of the environment and language instruction as input, then LLMs generate code that interacts with VLMs to produce a sequence of 3D affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation.
Table 1: Robot Foundation Models.This table organizes information about robot foundation models, specifically focusing on robotic manipulation tasks. In the ‘dataset’ section, only manipulation datasets are considered. Four collection methods are defined in ‘Collection Method’: Teleoperation (Human control a robot to perform tasks from a distance, such as the use of VR headsets with controllers), videos (Tasks performed by other robots or humans are recorded for the robot to mimic.), simulation (Tasks are collected in a simulation environment), and skill libraries (Pre-trained models are used to execute in new scene to enhance the dataset with the check module to recognize the task success). ‘Size’ refers to the parameters of the model. ”Input Modality” refers to the modality in which data is inputted into the model. ‘Architecture’ refers to the structure of the model. ‘Hardware’ refers to the devices used during the model training stage. ‘Output modality’ shares the same definition in Section 7. ’Frequency’ refers to the rate at which the model inference and control. ‘Benchmark’ refers to standard tests, focusing solely on manipulation, used to measure the performance.
Paper Robotics Dataset Collection Method Size Input Modality Architecture Hardware Output Modality Frequency Benchmark BC-Z (Jang et al. (2022)) BC-Z dataset (Jang et al. (2022)) Human Demonstrations1 Human Videos Robot Videos - Image Language or Video ResNet18 encoder + FiLM layers + a two-layer MLP 18 NVIDIA V100 GPUs Target Pose 10HZ Own benchmark (100 tasks in total) MOO (Stone et al. (2023)) RT-1 dataset (Brohan et al. (2022)) Self Collection Human Demonstrations  [^1] M Image Language RT-1 + OWL-ViT - Delta Pose - Own benchmark (1472 real world valuations) RT-1 (Brohan et al. (2022)) RT-1 dataset (Brohan et al. (2022)) Teleoperation 35M Image Language Conditioned EfficientNet + TokenLearnr + Decoder-only transformer - Delta Pose 3HZ Own benchmark (over 700 instructions) RT-2 (Brohan et al. (2023)) RT-1 dataset (Brohan et al. (2022)) Teleoperation 3/5/12/55B Image Language PaLI-X (Chen et al. (2022)), PaLM-E (Driess et al. (2023)) Multi-TPU Cloud Service Delta Pose 1-3HZ (Language Table Lynch et al. (2023)) Own benchmark RT-X (Padalkar et al. (2023a)) 60 individual datasets across 22 embodiments Human Demonstrations - Image Language PaLI-X (Chen et al. (2022)), PaLM-E (Driess et al. (2023)) Multi-TPU Cloud Service Delta Pose 3-10Hz Own benchmark (3600 evaluation trails) RT-H (Belkhale et al. (2024)) Diverse+Kitchen (Belkhale et al. (2024)) Human Demonstrations - Image Language PaLI-X (Chen et al. (2022)) - Delta Pose - Own benchmark (Kitchen tasks) LEO (Huang et al. (2023a)) CLIPort (Shridhar et al. (2021)) Human Demonstrations 1.3/7/13B Image 3D Point Cloud Language Decoder-only transformer 4 NVIDIA A100 GPUs Target Pose 1-5Hz Own benchmark (3 tasks for manipulation) RoboFlamingo (Li et al. (2023b)) CALVIN Bench(Mees et al. (2022)) Teleoperation 3/4/9B Image Language OpenFlamingo(Awadalla et al. (2023)) 8 NVIDIA Tesla A100 GPUs Delta Pose - CALVIN Bench(Mees et al. (2022)) Robocat (Bousmalis et al. (2023)) Meta-World (Yu et al. (2020)) DeepMind Control Suite (Tassa et al. (2018)) Self Collection Simulation Skill Library 364M/1.18B Image Language Proprioception Decoder-only transformer - Delta Pose Hindsight Image 10-20HZ RGB-Stacking Benchmark (Lee et al. (2021)) Roboagent (Bharadhwaj et al. (2023)) RoboSet(MT-ACT) (Bharadhwaj et al. (2023)) Teleoperation Skill Library - Image Language Proprioception Transformer One 2080Ti GPU Delta Pose 5Hz RL BENCH (James et al. (2020)) VIMA (Jiang et al. (2023)) VIMA dataset (Jiang et al. (2023)) Simulation 200M Image Language Decoder-only transformer 8 NVIDIA V100 GPUs Target pose - VIMA-BENCH (Jiang et al. (2023)) MIDAS (Li et al. (2023a)) VIMA dataset (Jiang et al. (2023)) Simulation 92M Image Language Decoder-only transformer 8 NVIDIA A100 GPUs Target pose - VIMA-BENCH (Jiang et al. (2023)) PaLM-E (Driess et al. (2023)) PaLM-E dataset (Driess et al. (2023)) Skill Library 12/22/562B Image Neural 3D Representations States of robot or objects Decoder-only transformer Multi-TPU Cloud Service Delta Pose Text 1-5Hz Language Table(Lynch et al. (2023)) Saycan(Ahn et al. (2022)) Gato (Reed et al. (2022)) Meta-World (Yu et al. (2020)) and other 15 datasets Simulation 1.2B Image Language Proprioception Decoder-only transformer TPU v3 Delta Pose 20HZ Meta-World (Yu et al. (2020)) GR-1 (Wu et al. (2023a)) Ego4D dataset (Grauman et al. (2022)) CALVIN Bench(Mees et al. (2022)) Self Collection Human Videos Teleoperation 195M Image Language Proprioception Decoder-only transformer - Delta Pose Image - CALVIN Bench(Mees et al. (2022)) QT-Opt (Kalashnikov et al. (2018)) QT-Opt dataset (Kalashnikov et al. (2018)) Simulation - Image Gripper Status Height Reinforcement Learning 10 NVIDIA P100 GPUs Target Pose - Own benchmark (28 test objects) RLS (Herzog et al. (2023b)) RLS dataset (Herzog et al. (2023b)) Skill Library (Simulation + Real World) - Image Proprioception CNN TPUv3 Target Pose - Own benchmark (9 scenarios) Q-Transformer (Chebotar et al. (2023b)) Q-Transformer dataset (Chebotar et al. (2023b)) Human Demonstrations Skill Library - Image Language Q-Transformer network - Target Pose 3HZ RT-1 (Brohan et al. (2022))
  • 1
    The articles mention the use of human demonstration but do not provide detailed descriptions of the specific methods employed.
    “-” donates no data is reported.
7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capabilities of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, a plethora of literature investigates the use of foundation models to hierarchically solve challenges like long-horizon problems and reward function design.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function.
Following Fig. 7 and Table. 1, pre-trained foundation models can engage with Policy modules in various forms. However, there is currently no benchmark indicating the most effective approach. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
8 Manipulation Data Generation
To propel robots into the era of general manipulation, the acquisition of vast amounts of data is indispensable (Padalkar et al. (2023b)). Gathering real-world data incurs substantial human labor costs. RoboVQA (Sermanet et al. (2023)) adopts a bottom-up data collection approach to gather specific data for fine-tuning VideoCoCa. This bottom-up data collection approach eliminates the need for human scene arrangement and fine-turned VideoCoCa can assist human to decide the next movement in data collection process. While collecting data in simulation presents a more economical solution, it still necessitates human effort in crafting both scene generation and task execution code tailored to specific tasks (Wang et al. (2023a)). The notorious sim-to-real gap issue persists in collecting data in simulation. Therefore, datasets for manipulating deformable and articulated objects still need to be acquired in the real world. Regardless of whether it’s in a real or simulated environment, improving the efficiency of data collection is essential. The mainstream approach is dataset augmentation, primarily divided into object-level and task-level augmentation.
Figure 8: Foundation Models for Manipulation Data Generation. 1) Object-Level Augmentation: The mask of the corresponding object is marked in the image through object segmentation, followed by generative rendering to modify the object’s visual or pose (Wang et al. (2023b)). 2) Task-Level Augmentation: Method 1 generates scene code and policy code by inputting new task instructions into LLM (Wang et al. (2023a)). In simulation scenarios, task execution success can be directly detected through check code, while in real-world scenarios, VLM can be used to detect task success (Ge et al. (2023)). In contrast to Method1, Method2 utilizes LLM to generate new object descriptions, followed by VGM to create new object meshes, forming corresponding scenes. However, the stability of code generation by LLM is comparatively lower. Hence, task instructions are decomposed into subgoals using LLM, followed by RL, Action Primitive, and Gradient Optimization algorithms to generate corresponding actions (Wang et al. (2023b)).
In object-level augmentation, the main idea is to use semantic segmentation to extract masks for each object, and then employ generative rendering methods to alter the object’s visual or pose (Mandlekar et al. (2023); Mandi et al. (2022)). GenAug (Chen et al. (2023c)) leverages language prompts with a generative model to modify object textures and shapes, adding new distractors and background scenes. ROSIE (Yu et al. (2023)) localizes the augmentation region with an open vocabulary segmentation model and then runs image editor to perform text-guided image editing.
In task-level augmentation, the main idea is to use LLMs to generate new task objectives or manually input new task objectives. VGMs is then used to generate meshes for the corresponding task scene objects, or LLMs directly generate simulation code for the corresponding task scene. VLMs or LLMs generate check code to determine the success of executing the new task. GenSim (Wang et al. (2023a)) leverages the code generation capability of LLMs to produce simulation task codes based on new task descriptions. RoboGen (Wang et al. (2023b)) surpasses GenSim in autonomy. It autonomously utilizes LLM to create task descriptions based on specific robot and object inputs and then uses VGM to generate 3D meshes based on these descriptions. Ha et al. (2023) decomposes the new task description into sub-goals using LLM, and then the plan is grounded into a sequence of 6 DOF exploration primitives. An LLM infers a success function code snippet and uses it to verify and label the task as succeeded or failed. PAFF (Ge et al. (2023)) notes that incorrect actions also correspond to new tasks and uses fine-tuned CLIP to label the erroneous trajectories with appropriate task labels. DIAL (Xiao et al. (2022a)) uses fine-tuned CLIP as automatic labelers for robot data, effectively importing internet-scale knowledge into existing datasets with limited ground truth annotations.
Following Fig. 8, most current foundation models for manipulation data generation focus on image rendering or simulation scenes. Exploring the use of foundation models in real-world scenarios to assist in real-time data labeling is also worthwhile.
9 Discussion
In this survey, we aim to outline the opportunities brought by foundation models for a general-purpose robot framework. We believe the potential of embedding foundation models into manipulation tasks as a viable path towards achieving general manipulation. However, the primary applications of LLMs, VFMs, VLMs, LMMs and VGMs focus only on certain aspects of general operational capabilities, such as reasoning, perception, multimodal understanding, and data generation. The current framework for RFMs resembles human learning, requiring no calibration or ontology parameters, directly learning the mapping between vision and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring safety remains an unresolved concern. Therefore, this paper proposes a framework of robot learning for manipulation towards achieving general manipulation capabilities and detailing how foundation models can address challenges in each module of the framework. However, there are still many open questions in this survey. In this section, we delve into several open questions that we are particularly concerned about.
9.1 What is the framework for general manipulation?
9.1.1 Definition of general manipulation.
The ultimate universal manipulation framework should be capable of interacting with humans to manipulate arbitrary objects in open-world scenarios and achieve diverse operational objectives. However, the interaction between agent and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios can be static or dynamic, with objects being either rigid or deformable, and task objectives ranging from short-horizon to long-horizon. We designate the restriction of the agent’s learning capability to improving old skills and manipulating rigid objects in static scenes to achieve short-horizon task objectives as Level 0 (L0), the current algorithm implementation has a high probability of achieving L0 product deployment. However, safety and accuracy remain paramount concerns.
9.1.2 The design logic of the framework in this survey.
From the definition and considerations of safety and accuracy at the L0 level, this paper proposes a framework for a general manipulation capability. Given that the scenarios are static, the framework is designed in a modular, sequential manner. To facilitate module migration, it is preferable for each module to be plug-and-play. Given the current reliance on human-in-the-loop mechanisms in autonomous driving and medical robotics to ensure safety, this framework aims for human-agent interaction through corrective instructions to ensure the safety of manipulation actions. The corrective action can be collected into the dataset and then improve old skills through offline training.
9.1.3 Product implementation strategy.
During agent operation, continuous human supervision is not always feasible. Hence, integrating real-time monitoring through parallel surveillance videos during robot execution could enhance safety. The framework in this paper does not explicitly denote this parallel safety monitoring module, as it resembles the post-conditions detection module. The post-conditions detection module analyzes the robot’s operation video to identify reasons for task failure, facilitating post-hoc correction to ensure task success. If the algorithm’s task execution safety is 80%, and the monitoring module predicts safety at 80% as well, the probability of risky movements reduces to 4%. Of course, for household robots, ensuring an over 99% safety rate is imperative. Initially, transitioning nanny from in-home caretakers to cloud-based monitoring of household robot operations appears to be the safest approach and a nanny oversees multiple household robots, allowing human intervention to promptly halt or correct hazardous behaviors. Later, by gathering extensive data to improve model accuracy. This framework still has shortcomings, such as the sequential execution order leads to inefficient characteristics.
9.2 How to design post-conditions detection and post-hoc correction?
The current data collection only focuses on gathering successful task execution data, ignoring the collection of data related to failed task executions. However, if data on failed task executions are collected and annotated with corresponding error reasons, it would be possible to train a model to both determine task execution success and analyze the reasons for task execution failure. For instance, RoboVQA (Sermanet et al. (2023)) utilizes crowdsourced data collection and trains videoCoCo (Yan et al. (2022)), employing a video-language model to analyze task execution description. Similarly, a post-conditions detection model could utilize a video-language model for training, taking manipulation videos as input to output whether tasks are successfully executed and to analyze the reasons for task execution failure. Post-hoc correction could then generate corrective action sequences based on the reasons for task execution failure and the task objectives, which would be handed over to a policy to generate corresponding corrective actions.
9.3 What kind of learning capability should a general manipulation framework possess?
9.3.1 The importance of learning ability.
As an intelligent agent for general manipulation, it is inevitable that one cannot learn all the skills of an open-world in one go, hence possessing a certain learning capability is necessary. Within the framework of this paper, a module of corrective instruction is introduced, enabling the robot to rectify its actions. These corrective demonstrations are incorporated into the manipulation dataset and used to improve the policy offline through fine-tuning. However, this approach still focuses on learning old task skills and cannot acquire new ones.
9.3.2 Definition of learning ability.
The model of Policy should possess the capability of interactive few-shot continue online learning to acquire a new skill and reinforce the policy’s mastery of the new skill through corrective instruction offline. Interactive refers to the ability to learn through human demonstration or by observing instructional videos. Learning through human demonstration often requires physical control or teleoperation, which is less natural. Learning through observation of instructional videos aligns better with human learning patterns. However, when humans learn from teachers, they often do not predict the teacher’s trajectory but rather understand the high-level description of the actions, akin to VLaMP (Patel et al. (2023)). Few-shot continue learning enables the agent to learn new skills with minimal demonstrations without forgetting previously learned skills. Online learning entails processing observed data instantly and enabling the model to learn as quickly as possible.
9.4 The strengths and weaknesses of current RFMs.
This paper categorizes models trained using manipulation datasets as RFMs. RFMs significantly enhance the policy’s generalization. For instance, LEO (Huang et al. (2023a)) directly outputs target pose, while RoboFlamingo (Li et al. (2023b)) directly outputs delta pose. Directly outputting target pose is easier to train compared to delta pose, although delta pose better aligns with human behavior. Currently, RFMs are fine-tuned using pre-trained models, but this deprives them of self-exploratory learning. Designing an architecture for RFMs is crucial. Utilizing foundation models to aid reinforcement learning still holds many unexplored areas, such as applying VLMs to dense reward function.
Training RFMs simultaneously requires a large amount of training data. Current data collection methods include gathering data in simulation, using skill library to add data, teleoperation, and imitation learning from human videos. While collecting data in simulation is cost-effective, it suffers from the sim-to-real gap. In order to add data using skill library, it is necessary to construct an initial skill library. Teleoperation can address the sim-to-real gap but comes with high hardware costs and the issue of unnatural third-person demonstration data. Imitation learning from human videos involves mapping human trajectories from videos to robot space, leading to issues with trajectory accuracy. This paper argues that the optimal data collection method involves individuals wearing exoskeleton devices equipped with various sensors. These devices accurately capture human actions and map them to robot operation space, enabling first-person data collection and addressing the sim-to-real gap.
9.5 How to establish a benchmark?
The challenge in establishing benchmarks for foundation models for manipulation lies in addressing diverse task objectives. Thus, a framework categorizing benchmarks for different tasks is necessary. This paper’s framework encompasses multiple tasks including interaction, hierarchical, perception, pre- and post-conditions detection, policy, and manipulation data generation. VLADP models cannot achieve transfer ability across different robot configurations, and the workspaces of different robot configurations also differ. Designing an evaluation function to incorporate robot configurations may alleviate this issue. Moreover, the real-time performance of foundation models also needs consideration.
10 Conclusion
In this survey, we aim to outline the opportunities brought by foundation models for a general-purpose robot framework. Except for RFMs, none of the foundation models have utilized manipulation task-level data to train, thus only capable of addressing a subset of functionalities for general operations. The current framework for RFMs resembles human learning, requiring no calibration or ontology parameters, directly learning the mapping between vision and action. However, this demands extensive data for learning, posing a crucial issue of constructing a data close-loop, and ensuring safety remains an unresolved concern.
As we propose a framework for a general-purpose manipulation agent, we first consider it to be language-conditioned, hence integrating an interaction module. The ability to understand corrective instruction within this module is crucial; agents inevitably make mistakes, and being able to correct them is pivotal on top of the three robotic fundamental principles. Understanding the environment and task execution in Pre- and Post-conditions Detection determines how tasks are completed and their measurement, while Post-hoc Correction serves as a check and self-correcting mechanism for robot safety and smooth task execution. Optimizing task execution plans within the Hierarchy of Skills is a decision-making process determining the ’intelligence’ of this agent, while skill execution decides the ’physical fitness’ of the agent.
Foundation models accelerate the research progress of general agents. LLMs and VFMs facilitate open-world interaction and perception, VLMs and LMMs ensure multimodal understanding. Current multimodal large models align modalities to text, as text tokens are conducive to reasoning, akin to humans contemplating information from books for cognitive processes. VGMs present significant opportunities for scene or object generation. However, each foundation model enhances the capability of zero-shot and few-shot learning. Autonomous driving is classified into levels L0-L5 . If we rank improving old skills and rigid static short-horizon operations as L0, the development of foundation models will bring excellent opportunities for general manipulation at L0.
While this survey raises some questions for consideration, the proposed framework for a general-purpose robot is sequential, whereas human behavior is often parallel. Nonetheless, we consider a sequentially connected framework as the first step towards L0 general-purpose manipulation agent. However, it’s undeniable that each module within the current framework faces numerous challenges:
  1. 1.
    Interaction In human interaction, there’s more than just language—it involves gestures or actions too. Adding multiple modal inputs to interaction modules can enhance recognition capabilities.
  2. 2.
    Hierarchical of skills The hierarchy of skills still has many unconsidered factors, such as achieving tasks in the shortest time with the highest efficiency, and how to devise proactive strategies for dynamic scenes.
  3. 3.
    Pre- and post-conditions detection Post-conditions detection needs to analyze reasons for task failure and enables self-correction within the post-hoc correction module. In current research, failure in task execution is typically determined by analyzing the status of objects after skill execution. But, post-conditions detection should instead analyze the robot’s own execution process to identify reasons for task failure. For instance, if a robot fails a task due to object collision during execution, post-hoc correction should prevent such occurrences during corrective actions. The perspective of observation might be a hidden challenge, requiring agents to actively perceive to analyze failure reason.
  4. 4.
    State The representation of state requires integration of multiple modalities, such as touch and hearing. At the same time, it’s important to consider the opportunities that foundation models can bring to active perception.
  5. 5.
    Policy The policy encompasses various RFMs, yet RFMs lack the capacity for self-exploration and a unified benchmark for comparing them, hindering RFMs development. Meanwhile, the extensive parameters of RFMs require significant computational resources for real-time operation, and model training also demands abundant data.
  6. 6.
    Environment Transition Module The foundation models inherently contain abundant physical priors. Applying foundation models to build a highly realistic physical model assist reinforcement learning training in transition models is a direction worth exploring.
  7. 7.
    Data Generation Output from LLMs and VGMs in data collection remains unstable, necessitating appropriate check module and data cleaning algorithms.